### v1

# EgoDTM: Towards 3D-Aware Egocentric Video-Language Pretraining

Abstract
Egocentric video-language pretraining has significantly advanced video representation learning. Humans perceive and interact with a fully 3D world, developing spatial awareness that extends beyond text-based understanding. However, most previous works learn from 1D text or 2D visual cues, such as bounding boxes, which inherently lack 3D understanding. To bridge this gap, we introduce EgoDTM, an Egocentric Depth- and Text-aware Model, jointly trained through large-scale 3D-aware video pretraining and video-text contrastive learning. EgoDTM incorporates a lightweight 3D-aware decoder to efficiently learn 3D-awareness from pseudo depth maps generated by depth estimation models. To further facilitate 3D-aware video pretraining, we enrich the original brief captions with hand-object visual cues by organically combining several foundation models. Extensive experiments demonstrate EgoDTM’s superior performance across diverse downstream tasks, highlighting its superior 3D-aware visual understanding. Our code will be released at https://github.com/xuboshen/EgoDTM.
1 Introduction
Figure 1: Comparison of different egocentric pretraining paradigms. While previous paradigms focus on text-based [33, 47, 78] or 2D spatial region-aware learning [72], EgoDTM incorporates 3D spatial information to enhance video representations.
The development of embodied AI capable of fulfilling diverse societal roles has long been a goal in artificial intelligence [23, 46]. A promising path to achieving this vision involves developing egocentric AI, which can comprehend human activities by analyzing large-scale egocentric videos captured from a first-person perspective using wearable devices. These videos provide rich insights into hand-object interactions (HOI) [18, 71, 10, 36], offering a window into how individuals interact with nearby objects through their hands. With the emergence of large-scale egocentric datasets [20], video-language pretraining [33, 78] has become a dominant paradigm for learning egocentric video representations, significantly improving performance on downstream tasks such as video-text retrieval [10, 60] and action recognition [59, 28].
Humans possess an innate ability to perceive and reason about 3D spatial relationships, effortlessly perceiving relative distances and spatial arrangements from visual cues alone [61, 25, 58]. While there is no universal definition of a model’s 3D awareness, we define it as the ability to infer 3D information from 2D images. However, models pretrained solely on 2D frames and general-purpose text (e.g., CLIP [49]) often struggle to develop robust 3D awareness [17, 39, 52]. To equip egocentric models with such 3D awareness, we wonder: Can video-language models better understand egocentric contexts by incorporating 3D-aware perception?
Achieving this goal presents two key challenges. First, 3D representations are diverse, and obtaining corresponding labels is often costly. Recent works [70, 76] explore using multi-view images for 3D reconstruction to distill 3D awareness into 2D models. However, in egocentric scenarios, depth maps are more practical and easier to obtain. Depth maps provide direct 3D distance information and help distinguish salient objects from the background, offering crucial cues for spatial understanding. Unfortunately, existing egocentric datasets with depth maps [36, 51, 26] remain limited in both scale and diversity for large-scale pretraining. Recent foundation models for depth estimation [68, 69, 5] demonstrate strong out-of-domain generalization, enabling us to construct a large-scale depth-augmented egocentric dataset at a low cost.
Second, effective 3D-aware pretraining requires bridging the modality gap between depth and text. As illustrated in Figure 1, common video-language pretraining [33, 47, 78] primarily relies on textual supervision, thus avoiding this challenge. Recent works [72, 43] explore region-aware pretraining using non-pixel-level cues, such as text and sparse object bounding boxes, which face only minor modality gaps. However, unlike these non-pixel-level cues, regressing the pixel-level outputs (e.g., depth maps) requires fundamentally different capabilities compared to predicting texts or bounding boxes, as suggested by previous studies [27, 31]. Additionally, depth estimation typically requires high-resolution inputs and multi-scale features, whereas video-language models rely on large batch sizes for contrastive learning, making direct integration nontrivial. Thus, designing an effective learning approach and enriching textual supervision with spatial information are crucial for successful 3D-aware pretraining.
To address these aforementioned challenges, we introduce EgoDTM, a novel 3D-aware egocentric video-language model that learns video representation from depth maps and spatially informed captions. In addition to dual transformer encoders for video-text alignment, EgoDTM adopts a 3D-aware module for video pretraining and a data construction pipeline to enrich captions with spatial information. To adapt the 3D-aware module for the video-language framework, we propose a unified visual representation with a lightweight depth decoder, supervised by depth predictions from foundation models [68]. Specifically, the lightweight depth decoder takes as input the visual representation from the last layer of the video encoder and predicts the composition of discrete adaptive bins [3, 4, 57, 32] to estimate low-resolution depth maps. Moreover, we adopt off-the-shelf visual foundation models [56, 54] to create high-quality HOI masks by first detecting HOIs, then selecting key frames, and finally tracking bidirectionally. To enrich the original captions with spatial information, we leverage a large language model (LLM) guided by the temporally consistent HOI masks. Through 3D-aware video-language pretraining, EgoDTM improves visual generalization on downstream tasks involving egocentric HOIs.
Our contributions are threefold: (1) We introduce EgoDTM, a 3D-aware egocentric video-language model learned from 3D-aware video-language pretraining. (2) We develop a lightweight 3D-aware decoder for depth estimation and a data construction pipeline to enrich captions with spatial information. As a byproduct, we generate millions of egocentric data, including captions, HOI boxes, HOI masks, and depth maps. (3) Extensive experiments demonstrate that EgoDTM significantly enhances performance on video understanding tasks like video-text matching, and 3D understanding tasks like robot manipulation.
2 Related Works
Egocentric Video-Language Pretraining. Egocentric video understanding [20, 7, 45, 75, 40, 66, 10, 64, 67, 29, 65] typically involves human daily interaction between hands and objects from a first-person perspective. Inspired by visual-language pretraining paradigms in third-person datasets [2, 38, 49, 43, 72], EgoVLP [33] firstly proposes to conduct egocentric video-language pretraining on the large-scale egocentric dataset Ego4D [20], which aims to learn video representations from massive video-text data via contrastive learning. Since EgoVLP, similar paradigms [33, 47, 78, 1] have gained large success in EgoHOI understanding, For example, LaViLa [78] employs a video-conditioned GPT-2 [48] and a T5 [50] rephraser to generate text descriptions for egocentric videos to expand video-text data. Despite the progress, pretraining with text alone often lacks precision in target localization. In response, some studies have focused on developing region-aware representations for EgoVLMs. For example, HelpingHands [72] proposes learning from noisy hand-object detection results, while HENASY [43] ensembles an additional hierarchical encoder to learn HOI region-aware representation. However, these methods are limited to 2D reasoning and lack an understanding of real-world 3D context. In this work, we take a step towards 3D-aware egocentric video-language pretraining.
Figure 2: EgoDTM learns 3D-aware representations from depth and text. Our dual encoders are constructed using only transformers [15, 63, 49] with flash attention [11]. During pretraining, we conduct (1) 3D-aware video pretraining: we design a lightweight 3D-aware decoder to predict depth using visual feature maps, supervised by a teacher foundation model [69]. The decoder contains a plain feature pyramid to get multi-scale features, a depth-aware transformer decoder to process depth queries with video features, and the heads to predict depth maps; (2) Spatial-aware textual enrichment: we enhance captions with spatial information by organically combining foundation models in the detect-track-generate pipeline. Different green markers denote inconsistency of HOI predictions; identical ones indicate consistency.
3D-Awareness of 2D Vision Models. As suggested by research in developmental psychology and psychophysics, we humans are capable of understanding 3D information, like depth and orientations, from only 2D visual signals [61, 25, 58]. With the growing interest in embodied intelligence [23, 41] and personal AI assistants [46], the demand for 3D-aware 2D visual models has become increasingly important. Although there is no widely accepted consensus to define and build effective 3D-aware models, some studies have started to assess the 3D awareness of visual foundation models. For example, Probing3D [17] demonstrated that CLIP, which is pretrained exclusively with text, is significantly limited in predicting depth and 3D surface normals. EysWideShut[62] revealed that CLIP struggles with 3D-related tasks, such as recognizing object orientation and understanding compositional contexts. More recently, relevant works have made endeavors to enable visual foundation models (e.g., CLIP) with 3D awareness by fine-tuning on 3D-aware data. For instance, FiT3D [70] finetunes visual foundation models with 3D Gaussian features reconstructed from multi-view images. SpatialVLM [8] conducts visual instruction tuning on metric depth-aware QA pairs for multi-modal large language models. Different from these learning paradigms, we define 3D-awareness as the model’s latent ability to estimate depth and pretrain an egocentric video-language model with a lightweight 3D-aware decoder.
3 Method
Our objective is to develop a 3D-aware egocentric video-language model, where 3D-awareness is defined as the model’s ability to infer depth information from its representations. EgoDTM achieves this by pretraining depth-aware and text-aligned video representations.  Figure 2 illustrates the architecture of our proposed EgoDTM. We first describe the video-language model architecture in Section 3.1. Next, we introduce the 3D-aware video pretraining in Section 3.2 and then present our approach to generate spatial-aware captions in Section 3.3. The details of training and inference are presented in Section 3.4.
3.1 Video-Language Model Architecture
The video-language model follows a dual-encoder design [2], consisting of a video encoder and a text encoder.
Video and text encoders. Our video encoder employs a plain vision transformer [15] backbone to process video tokens. The video input $V\in\mathbb{R}^{F\times H\times W\times 3}$ is divided into $N=\frac{F}{F_{0}}\times\frac{H}{H_{0}}\times\frac{W}{W_{0}}$ non-overlapping cubes of dimension $F_{0}\times H_{0}\times W_{0}\times 3$, where $F$ is the number of frames, $H$ and $W$ denote height and width. These cubes are combined with positional embeddings and processed by the video encoder to produce a feature map $\mathbf{v}_{Z}\in\mathbb{R}^{N\times C}$. Note that a ${\rm[CLS]}$ token is added to represent the global video embedding $\mathbf{v}_{\rm cls}$. For the text encoder, captions $T$ are tokenized via a Byte Pair Encoding (BPE) tokenizer [55] and encoded with a transformer initialized from CLIP [49] to output the sentence embedding $\mathbf{t}_{\rm cls}$. Both encoders adopt the flash-attention [11] to mitigate the memory bottleneck of the attention mechanism.
Video-text alignment loss. We use the InfoNCE loss to align video and text embeddings. For simplicity, we define the video-to-text loss within a batch =$\mathcal{B}=\{(\mathbf{v}_{i},\mathbf{t}_{i}\}_{i=1}^{B}$ as follows:
$\mathcal{L}_{\rm vtc}=-\frac{1}{B}\sum_{i}{\rm log}\frac{{\rm exp}(\mathbf{v}_{i}\cdot\mathbf{t}_{i}/\tau)}{\Sigma_{j}{\rm exp}(\mathbf{v}_{i}\cdot\mathbf{t}_{j}/\tau)}$ (1)
where $(\mathbf{v}_{i},\mathbf{t}_{i})$ represent the video and text embeddings $(\mathbf{v}_{\rm cls},\mathbf{t}_{\rm cls})$ for the $i$-th (video, caption) pair within the batch. The text-to-video loss is defined symmetrically.
3.2 3D-Aware Video Pretraining
A critical issue in leveraging depth maps for 3D-aware pretraining is that depth maps are typically high-resolution images (e.g., 1024p). However, most low-level details in these maps are redundant, as video-language models primarily benefit from the relative depth information between patches. Therefore, high-resolution depth maps may be unnecessary and inefficient for pretraining. Additionally, video-text pretraining often requires a large batch size to ensure good performance, further emphasizing the need for lightweight and memory-efficient processing. To address these concerns, we design a lightweight 3D-aware decoder to estimate low-resolution depth maps from video representations. Specifically, the decoder consists of a plain feature pyramid, a depth-aware transformer decoder, and two specialized heads, as described below.
Plain feature pyramid. Depth estimation methods [44, 3, 53] often utilize multi-scale features from intermediate backbone layers. However, video-language pretraining typically uses only the final-layer representations. To reconcile these differences, we adopt a simplified FPN [30] that projects the video feature map $\mathbf{v}_{Z}$ into multi-scale outputs $\{\mathbf{v}_{Z_{1\times}},\mathbf{v}_{Z_{2\times}},\cdots,\mathbf{v}_{Z_{m\times}}\}$, where $\mathbf{v}_{Z_{m\times}}$ denotes the feature map of size $(N\times m^{2})\times C$, and the size of the predicted depth maps equals $N\times m^{2}$. Moreover, inspired by the adaptive bins [3, 4, 57, 32] in depth estimation, we segment the depth range into intervals and predict distributions at each pixel over the interval centers.
Depth-aware transformer decoder. This module processes multi-scale features to generate depth-aware representations capable of estimating distinct distance ranges. The transformer decoder follows standard architecture as [6, 9], transforming $L$ learnable depth queries $\mathbf{Q}\in\mathbb{R}^{L\times C}$ into depth-aware representations $\mathbf{Z}\in\mathbb{R}^{L\times C}$ using self-attention and cross-attention mechanisms. The cross attention processes $\mathbf{Q}$ as queries and concatenates $(\mathbf{v}_{Z_{1\times}},\dots,\mathbf{v}_{Z_{m-1\times}})$ as key-value pairs $(k,v)$. The learned depth queries estimate depths at different distances, as illustrated by varied gray colors in Figure 2.
Bin head. We partition the depth range of $[0,1]$ into $L$ intervals. The bin head $\mathcal{F}(\cdot):\mathbb{R}^{C}\rightarrow[0,1]$ processes depth-aware representations $\{\mathbf{z}_{i}\}_{i=1}^{L}$ to predict widths of depth interval bins $\{w_{i}\}_{i=0}^{L}$, satisfying $\sum_{i=0}^{L}w_{i}=1$ where we denote $w_{0}=0$. The $i$-th depth interval is defined by boundaries $[\sum_{i=1}^{i-1}w_{i},\sum_{i=1}^{i}w_{i}]$, with its center computed as:
$c_{i}=\sum_{i=1}^{i-1}w_{i}+\frac{w_{i}}{2}$ (2)
Mask head. Then, we aim to predict distributions over the depth centers at each pixel. Specifically, the mask head $\mathcal{G}(\cdot):\mathbb{R}^{C}\rightarrow\mathbb{R}^{C}$ takes $\mathbf{z}_{i}$ as input, and the output $\mathbf{m}_{i}$ is multiplied by the largest feature map $\mathbf{m}_{i}\otimes\mathbf{v}_{Z_{m\times}}$ to acquire unnormalized probabilistic maps $d_{i}\in\mathbb{R}^{N\times m^{2}}$, where $\otimes$ denotes the composition of element-wise multiplication and sum along the feature dimension. After normalizing $\{d_{i}\}_{i=1}^{L}$ by the Softmax operation along the query dimension $L$, we get the probabilistic map $D=[D_{1},\dots,D_{L}]\in\mathbb{R}^{L\times N\times m^{2}}$. The process is described as:
$D=\mathrm{Softmax}(\mathcal{G}(\mathbf{Z})\otimes\mathbf{v}_{z_{m\times}})$ (3)
Finally, depth prediction $\tilde{D}\in\mathbb{R}^{N\times m^{2}}$ is obtained by the linear combination of the depth-bin-centers and the probabilistic maps:
$\tilde{D}=\sum_{k=1}^{L}c_{k}D_{k}$ (4)
3D-aware pretraining loss. Given the lack of depth annotations in large-scale egocentric videos [20], we generate pseudo-depth labels using monocular depth estimation foundation models [53, 68, 69]. The pretraining loss is defined as:
$\mathcal{L}_{\rm depth}=\|\tilde{D}-D_{{\rm gt}}\|_{2}$ (5)
where $\tilde{D}$ and $D_{\rm gt}$ are the predicted and ground-truths inverse depth maps within the range $[0,1]$, respectively.
3.3 Spatial-Aware Textual Enrichment
Previous egocentric video-language pretraining frameworks typically learn from short descriptions composed of verbs and nouns. To foster 3D-aware pretraining, we argue that captions should contain more spatial-temporal information like HOI positions, object shapes, and movements. We therefore enrich the captions with visual cues by employing a detect-track-generate pipeline tailored for egocentric videos using foundation models [56, 54, 13].
Detect. We use the detector [56], HOID, finetuned on egocentric data to detect hands, objects, and their contact states for a given frame. The detector $f(\cdot)$ receives $V_{i}$ as input and outputs diverse information $\{(b_{j},c_{j},t,s)\}_{j}$, where $b_{j}$ is the bounding boxes with confidence scores $c_{j}$, $t$ is object type, and $s$ is contact state of hand.
Track. Since predictions of HOID lack temporal consistency, we address this issue by tracking across the video. We utilize SAM2 [54] $g(\cdot)$, a foundation model for promptable segmentation and tracking. The initial tracking frame $V^{*}$ is selected when the hands make contact with objects. The HOI masks of this frame are obtained by prompting SAM2 with HOI boxes $\{m^{*}_{j}\}_{j}=g(V^{*},\{b^{*}_{j}\}_{j})$, where the bounding boxes serve as negative prompts for each other. Finally, the model tracks HOI masks throughout all frames to acquire the temporal consistent masks $M=(\{m_{1j}\}_{j},\dots,\{m_{Fj}\}_{j})$.
Generate. As suggested by previous works [34, 19], LLMs are capable of understanding visual cues like bounding boxes. We use an LLM $h(\cdot)$ to comprehend the HOI spatial information and original captions, then generate spatial-aware captions $T^{\text{aug}}=h(T,M)$.
Figure 3: Qualitative results of depth estimation from EgoDTM and DepthAnythingv2 [69](DAv2) on datasets including in-domain but unseen Ego4D validation set [20], out-of-domain and unseen data of EK100 [10], MECCANO [51], and H2O [26]. Note that DAv2 operates with a high resolution of 512p, while EgoDTM uses a lower resolution input of 224p and generates a depth map at a resolution of 56p. Despite the lower resolution input, EgoDTM demonstrates intuitive generalization across unseen egocentric datasets with diverse environments, illuminations, backgrounds, and varying HOI object sizes.
Equipped with the enriched video-text pairs, our video-text contrastive learning objective is represented as:
$\mathcal{L}_{\rm augvtc}=-\frac{1}{B}\sum_{i}{\rm log}\frac{{\rm exp}(\mathbf{v}_{i}\cdot\mathbf{t}_{i}^{\prime})/\tau)}{\Sigma_{j}{\rm exp}(\mathbf{v}_{i}\cdot\mathbf{t}_{j}^{\prime}/\tau)}$ (6)
where $\mathbf{t}_{i}^{\prime}={\rm rand}(\mathbf{t}_{i},\mathbf{t}_{i}^{\rm aug})$ is randomly sampled from the original and enriched text embedding $\mathbf{t}_{i}^{\rm aug}$.
3.4 Training Strategy
Training and evaluation. We employ both the augmented video-text alignment loss and the 3D-aware pretraining loss to pretrain our model:
$\mathcal{L}=\mathcal{L}_{\rm augvtc}+\mathcal{L}_{\rm depth}$ (7)
Our pretraining improves video-language representations for downstream egocentric tasks without modifying the dual-encoders themselves. As a result, only the dual-encoders are used for downstream tasks, introducing no additional computational costs at inference time.
Pretraining data. Our pretraining data consists of four million (video, text) pairs, with each video approximately 1 second long. We generate enhanced text descriptions and depth maps specifically for videos featuring hand-object interactions, resulting in two million (video, enriched texts, depths) triplets. Specifically, for each video, we extract eight depth maps using the DAv2-Large [69] model. The HOI boxes are detected by HOID [56], which is finetuned on egocentric data. The HOI masks are segmented and tracked using the same unified SAM2-Large model [54]. The final enriched text descriptions are generated using the DeepSeek-LLM-200B model [13].
Epic-Kitchens-100-MIR EGTEA EgoMCQ
Method Supervision Text Param mAP (%) nDCG (%) mean-acc top1-acc Inter-Acc Intra-Acc
[…]
EgoVLP [33] Text 4M 172M 26.0 20.6 23.3 28.8 27.0 27.9 - - 90.6 57.2
EgoVLPv2 [47] Text 4M 363M 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 91.0 60.9
LaViLa [78] Text 35M 169M 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 93.6 59.1
AVION [77] Text 35M 142M 37.1 28.7 32.9 34.4 31.0 32.7 38.6 42.3 94.4 62.1
HelpingHands [72] Text, Box 35M 206M 35.6 26.8 31.2 34.7 31.7 33.2 29.4 35.3 93.2 58.8
HENASY [43] Text, Box 35M 277M 35.5 27.1 31.3 34.6 31.7 33.2 29.6 35.9 94.1 61.3
EgoDTM (ours) Text, Depth 6M 151M 37.9 29.1 33.5 34.8 31.9 33.4 40.2 43.2 94.6 63.6
Table 1: Comparison with state-of-the-art methods on zero-shot video-text retrieval and action recognition. All methods use the same 4M videos from Ego4D for pretraining and have similar hidden dimension sizes. The numbers of the method with are sourced from [43].
4 Experiments
To evaluate our EgoDTM, we conduct experiments from three perspectives: short video understanding (video-text retrieval and action recognition), 3D space understanding (depth estimation and robot manipulation), and long video understanding (natural language query). These evaluations cover seven benchmarks across five datasets. In the following sections, we detail the experimental setups (Section 4.1), main results (Section 4.2), ablation studies (Section 4.3), and further analyses (Section 4.4).
4.1 Benchmarks and Settings
Zero-shot video-text retrieval (ZS-VTR). To assess video-text alignment, we conduct zero-shot retrieval evaluations on text-to-video multi-choice retrieval on Ego4D [20] (EgoMCQ) and multi-instance video-text retrieval on Epic-Kitchens-100 [10] (EK-100-MIR). Following [33, 10], we use mean Average Precision (mAP) and the normalized Discounted Cumulative Gain (nDCG) as metrics for EK-100-MIR and accuracy for EgoMCQ.
Zero-shot action recognition (ZS-AR). Action recognition is conducted in a video-to-text retrieval manner. We experiment on the EGTEA [28] dataset, requiring models to recognize 106 classes of cooking activities. The metrics include mean accuracy and top-1 accuracy across test splits.
Depth estimation (DE). To assess the model’s 3D awareness, we conduct depth estimation by finetuning our model on the H2O [26] dataset. Following [17], we add a prediction head composed of a linear layer and an upsampling convolution layer on top of the frozen video encoder. Similar to [17, 16], the metrics are threshold accuracy ($\delta_{i}$): percentage of pixels $\tilde{D}_{j}$ in ground-truth depth maps that satisfy $\text{max}(\frac{D_{{\rm gt}j}}{\tilde{D}_{j}},\frac{\tilde{D}_{j}}{D_{{\rm gt}\ j}})=\delta_{i}<thr^{i}$ for $thr=1.25,i\in\{1,2,3\}$, and RMSE: $\|\tilde{D}-{D}_{\rm gt}\|_{2}$. The scale-invariant metrics are shifted and normalized from scale-aware metrics.
Method Scale-Aware Metrics Scale-Invariant Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE↓ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE↓
ConvNext [37] 0.721 0.965 0.991 0.644 0.727 0.969 0.996 0.593
CLIP [49] 0.795 0.966 0.988 0.624 0.811 0.976 0.994 0.559
EgoVLP [33] 0.778 0.954 0.989 0.610 0.853 0.977 0.996 0.497
LaViLa [78] 0.801 0.954 0.987 0.598 0.811 0.964 0.993 0.552
AVION [77] 0.786 0.960 0.991 0.606 0.812 0.969 0.996 0.543
EgoDTM (ours) 0.826 0.964 0.993 0.539 0.848 0.977 0.998 0.481
Table 2: Comparisons of depth estimation task on H2O dataset.
Robot manipulation (RM). The robot must learn to accurately perceive and interact in 3D spaces to accomplish daily tasks. Following previous works [42, 24], we evaluate the visual representations as frozen perception modules for downstream policy learning within the Franka Kitchens simulation environment [21]. Five tasks are adopted: turn knob (TK), open door (OD), flip switch (FS), open microwave (OM), and slide door (SD). We use success rate as the evaluation metric.
Natural language queries (NLQ). The NLQ task localizes time intervals in a long video given a language query. We experiment on the EgoNLQ task on Ego4D [20]. To fairly evaluate different pretrained visual representations, we extract the visual features by video encoders and text features by the same pretrained BERT [14], then train a VSLNet [73] to solve this task. The evaluation metrics are “Rn@m”, where $n\in\{1,5\}$ and $m\in\{0.3,0.5\}$, presenting the percentage of at least one of the top-$n$ predicted intervals having IoU greater than $m$.
Pretraining details. The dual-encoders are initialized by the checkpoint pretrained on the four million video-text pairs from Ego4D. EgoDTM is then trained for two epochs on 8*A800 GPUs, which requires approximately 10 hours and a learning rate of 3e-5. The depth-aware transformer decoder comprises six layers. The hidden dimension of the dual encoders is 768, while the 3D-aware decoder uses a dimension of 256 for efficient design. We use frames with 224p as input and 56p as output of the depth maps. Consequently, our 3D-aware decoder only has 9M parameters, and the batch size is set to 4096.
Method TK OD OM FS SD Average
R3M [42] 53.3% 50.7% 59.3% 86.3% 97.7% 69.4%
MPI [24] 83.3% 54% 44.5% 93.5% 100% 75%
ResNet [22] 28% 18% 26.7% 50% 75.5% 39.7%
CLIP [49] 26.3% 13% 24.7% 41.7% 86.3% 38.4%
LaViLa [78] 48% 26% 22.5% 69% 94.5% 52%
EgoDTM (ours) 56% 28% 35.5% 81% 92.5% 58.6%
Table 3: Comparison of robot manipulation tasks in Franka Kitchen simulation to assess model’s 3D-awareness.
4.2 Main Results
EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
$\mathcal{L}_{\rm vtc}$ 29.7 30.7 94.2 60.2
$\mathcal{L}_{\rm vtc}+\mathcal{L}_{\rm depth}$ 31.3 31.2 94.2 62.6
$\mathcal{L}_{\rm augvtc}$ 31.3 32.2 94.0 61.6
$\mathcal{L}_{\rm augvtc}+\mathcal{L}_{\rm depth}$ 33.1 33.1 94.6 62.6
(a) EgoDTM’s components. Both the 3D-aware video pretraining and textual enhancement enhance video-language representations.
EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
4 32.6 32.5 94.61 62.61
8 33.1 33.1 94.6 62.64
16 32.3 32.4 94.72 62.46
(b) Depth query numbers. We set 8 queries as the default, which empirically generalizes best on EK100MIR.
Reso Mem EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
28p 10G 30.4 31.2 94.54 60.65
56p 17G 30.9 31.9 94.28 61.05
112p 59G 31.2 31.9 94.37 60.38
(c) Resolution of depth. The batch size is set to 512 to accommodate the high GPU memory demands of large depth maps.
EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
Base 33.1 33.1 94.6 62.64
Large 36.4 34.2 95.13 66.04
(d) Model size. Larger model sizes bring more performance gain, especially on intra accuracy of EgoMCQ.
EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
All 32.8 32.7 94.46 61.92
Rand 33.1 33.1 94.6 62.64
(e) Text Sampling Strategy. Random substitution is better than thorough replacement.
EK-100-MIR EgoMCQ
mAP nDCG Inter Intra
1024 32.4 32.5 94.55 61.48
2048 33.1 33.1 94.6 62.64
4096 33.5 33.4 94.54 63.55
(f) Batch Size. Increasing the batch size steadily yields larger improvements.
Table 4: Zero-shot ablation studies. Models are pretrained on 4M video-text pairs from Ego4D and evaluated with zero-shot video-text retrieval on Epic-Kitchens-100-MIR [10] and EgoMCQ [33]. Unless specified, the default setting includes: ViT-B/16 backbone, batch size of 2048, random substitution of enriched texts, both $\mathcal{L}_{\rm depth}$ and $\mathcal{L}_{\rm augvtc}$ as losses, 8 depth queries, and 56p depth maps.
Zero-shot video-text retrieval. In Table 1, EgoDTM outperforms models like LaViLa [78] and AVION [77]. For example, on EK100-MIR, EgoDTM outperforms the state-of-the-art AVION model by +0.9% in mAP and +0.3% in nDCG, while also achieving a +1.5% improvement in intra accuracy on EgoMCQ. This indicates that captions generated based on visual cues from our data curation pipeline are more accurate and informative than captions generated from visual-conditioned LLMs or simply text rewriting. Besides, our model surpasses HelpingHands [72] and HENASY [43], which rely on noisy HOI detection supervision with extra parameters in the visual encoder, indicating that depth modality can potentially offer larger merits for video-language models.
Zero-shot action recognition  Table 1 illustrates the superior performance of EgoDTM on EGTEA. While previous works have demonstrated that depth modality enhances traditional action recognition tasks [74, 35], their applicability has been largely constrained to specific scenarios and closed-set datasets. Our results extend this understanding, showing that video-language models can effectively leverage depth modality to learn more generalizable video representations.
Depth estimation. We evaluate the ability of visual-language models to infer depth from images in Table 2. Our model surpasses all previous video-language models across most metrics by a large margin. In particular, our method improves the scale-aware RMSE by approximately 9.8% over the second-best result achieved by LaViLa. Interestingly, EgoVLP performs better in scale-invariant $\delta_{1}$ metric, suggesting that it latently captures finer-grained pixel-level details. We hypothesize that this advantage stems from EgoVLP’s scene-aware negative sampling strategy, which samples video-text pairs captured from the same environments into the same batch, thereby enhancing its ability to learn specific details.
Robot manipulation. As shown in Table 3, EgoDTM consistently outperforms pretrained visual-language models such as CLIP [49] and LaViLa [78] by +20.2% and +6.6%, respectively, demonstrating stronger spatial perception in visual representations. Additionally, EgoDTM performs competitively with specialized robot learning methods on certain tasks, such as “turn knob” but underperforms on others, like “open microwave”. A possible reason is that our model, pretained on depth and text, may overfit to real-worold scenarios, whereas methods like R3M [42] leverage self-supervised pretraining, which could provide better generalization to diverse manipulation tasks.
Method R1@0.3 R5@0.3 R1@0.5 R5@0.5
EgoVLP [33] 6.32 13.84 3.41 8.80
LaViLa [78] 7.12 14.82 3.87 9.55
AVION [77] 7.33 14.89 4.31 9.53
EgoDTM (ours) 8.13 16.11 4.83 10.30
Table 5: Comparisons of NLQ task on Ego4D dataset.
Natural language queries. As shown in  Table 5, EgoDTM outperforms other egocentric video-language models, e.g., +1.22 on R5@0.3 over AVION. In long video localization tasks, depth awareness enhances spatial understanding, enabling EgoDTM to better comprehend visual content.
4.3 Ablation Study
Ablation results are shown in Table 4.
EgoDTM’s components. Combining video-text matching with depth estimation enhances each learning objective. EgoDTM leverages depth maps to capture object relations, while textual descriptions enriched with shape and movement details can benefit more from depth information.
Depth query numbers. The number of depth queries affects depth granularity. We find that setting the query number to 8 is optimal, with each query covering a moderate depth range within $[0,1]$.
Resolution of the predicted depth map. The resolution of depth maps greatly impacts GPU memory usage. We find that higher-resolution depth maps slightly improve the video-text alignment, but this comes at the cost of reducing the maximum batch size. Therefore, we choose a resolution to 56p as a balanced trade-off between high resolution and the ability to maintain a large batch size.
Model size. A larger model size enhances performance, especially on the EgoMCQ intra task, which requires selecting the correct video from visually similar alternatives.
Text sampling strategy. We examine the impact of text sampling strategy on model performance. A mixed sampling strategy that includes both enriched and original texts for pretraining yields stronger results by leveraging both detailed and simplified textual information.
Batch size. The experiment results show that increasing batch size steadily improves performance.
4.4 More Analysis
Method mAP @(0.5:0.95) map @0.50 hand-AP @(0.5:0.95) object-AP @(0.5:0.95)
Upperbound [12] 60.70 73.31 90.87 30.53
Our Pipeline 43.02 54.88 61.01 25.03
Table 6: Evaluation of our HOI mask generation pipeline on HOI segmentation benchmark VISOR [12] in kitchen environment.
Modality mean-acc top1-acc top5-acc
RGB 61.6 68.2 87.4
RGB+Depth 62.5 69.5 89.2
Table 7: Multi-modal action recognition on EGTEA using our predicted depths. We apply a simple MLP to encode depth maps and perform multi-modal feature fusion via late fusion.
Figure 4: Qualitative results of robot manipulations.
Quantitative and qualitative analyses of the HOI mask generation pipeline. To evaluate the reliability of the HOI detector and SAM2 model in the egocentric domain, particularly for hand-object interactions, we evaluate the image-based HOI segmentation on the VISOR dataset [12]. The results are presented in Table 6. The supervised training model on VISOR serves as the upper bound for performance comparison. While our generation pipeline does not fully reach the upper bound, it demonstrates high-quality segmentation results. Notably, hand segmentation is significantly lower than the upper bound, likely due to inaccuracies in HOI detection prompts and challenges in segmenting hand-object interactions with SAM2 in cluttered egocentric backgrounds.  Figure 5 presents that our data construction pipeline is useful even when HOI regions are small while the background is noisy on Ego4D.
Visualization of predicted depth. In Figure 3, we visualize predicted depth maps in unseen environments. Despite operating with low-resolution inputs, EgoDTM effectively estimates depth, yielding results close to those of DAv2. Notably, there is a clear correlation between the depth maps and the hand-object interaction regions.
Predicted depth contains valuable information for action recognition. Since ground-truth data is unavailable for directly evaluating the depth decoder, we demonstrate the utility of our predicted depth maps for multimodal action recognition, as shown in  Table 7. We simply encode the depth map using an MLP at a 56p resolution. The action recognition accuracy improved by +1.4%, confirming that our predicted depths contain meaningful information for unseen egocentric data.
Case Study on Robot Manipulation. In Figure 4, the learned policy based on EgoDTM visual representation enables the robot to approach the switch and turn it, while LaViLa successfully approaches but misses the switch.
Figure 5: Example of generalizable data construction. For better visualizations, we blur the background to highlight HOI regions.
5 Conclusion
In this work, we present EgoDTM, a novel egocentric 3D-aware video-language model. EgoDTM integrates dual transformer encoders with a lightweight 3D-aware depth decoder, trained using video-text contrastive learning and depth estimation objectives. To enable large-scale pretraining, we generate millions of depth maps and spatially enriched captions by leveraging foundation models. The captions are enhanced through a detect-track-generate pipeline specifically tailored for egocentric videos. EgoDTM demonstrates intuitive generalization in estimating depths in unseen environments. Extensive experiments across diverse benchmarks—spanning short video understanding, 3D understanding, and long video understanding—validate the effectiveness of our approach.
Limitations. While EgoDTM achieves strong performance in egocentric scenarios, its impact on broader domains remains limited. We will explore pretraining 3D-aware video representations for general domains in the future.